OpenStudy (anonymous):

lim h->0 1/(sqrt x+h+2) +(sqrt x+2) find f'(x)

OpenStudy (anonymous):

looks like you are one step away

OpenStudy (anonymous):

lim goes to 1/sqrt(x+2)+sqrt(x+2), I suspect you mistyped something

OpenStudy (anonymous):

replace h by 0 and you get \[\frac{1}{2\sqrt{x+2}}\]

OpenStudy (anonymous):

which is another way of saying that if \[f(x)=\sqrt{x+2}\] then \[f'(x)=\frac{1}{2\sqrt{x+2}}\]

OpenStudy (anonymous):

which i am guessing is the original question

OpenStudy (anonymous):

kk great thx